OpenStudy (anonymous):

what is the recursive formula for the sequence... 15,26,48,92..

OpenStudy (amistre64):

well, we can start by seeing if it linear, quadratic or other by finding the difference between terms until they are alike

OpenStudy (anonymous):

f(n+1)=f(n)*2-4

OpenStudy (amistre64):

92 48 26 -48 -26 -15 ---- ---- ----- ; it aint linear 44 22 11 44 22 -22 -11 ---- ---- ; and it aint quadratic 22 11

OpenStudy (amistre64):

it might be exponential or logistic; but i aint got that far ....

OpenStudy (amistre64):

Ente might have it; i havent tried :)

OpenStudy (anonymous):

lets see... f(0)=15 f(1)=15*2-4=30-4=26 check f(2)=26*2-4=52-4=48 check f(3)=48*2-4=96-4=92 check so f(n+1)=f(n)*2-4 is a correct recursive formula for the problem

OpenStudy (amistre64):

yay!! :)

OpenStudy (anonymous):

correct ente

OpenStudy (amistre64):

either you have a good eye for it; or you know something more....

OpenStudy (anonymous):

just tried around a bit... loved that sort of question in school ^^ often harder to get an explicit formula though ;)